### v1

# V2I-Calib: A Novel Calibration Approach for Collaborative Vehicle and Infrastructure LiDAR Systems

Abstract
Cooperative vehicle and infrastructure LiDAR systems hold great potential, yet their implementation faces numerous challenges. Calibration of LiDAR systems across heterogeneous vehicle and infrastructure endpoints is a critical step to ensure the accuracy and consistency of perception system data, necessitating calibration methods that are real-time and stable. To this end, this paper introduces a novel calibration method for cooperative vehicle and road infrastructure LiDAR systems, which exploits spatial association information between detection boxes. The method centers around a novel Overall IoU metric that reflects the correlation of targets between vehicle and infrastructure, enabling real-time monitoring of calibration results. We search for common matching boxes between vehicle and infrastructure nodes by constructing an affinity matrix. Subsequently, these matching boxes undergo extrinsic parameter computation and optimization. Comparative and ablation experiments on the DAIR-V2X dataset confirm the superiority of our method. To better reflect the differences in calibration results, we have categorized the calibration tasks on the DAIR-V2X dataset based on their level of difficulty, enriching the dataset’s utility for future research. Our project is available at https://github.com/MassimoQu/v2i-calib .
I INTRODUCTION
Perception systems are crucial for ensuring the safe and effective operation of autonomous vehicles [1]. Vehicle-to-infrastructure (V2I) cooperative systems can enhance the reliability of perception systems in complex traffic scenarios or adverse weather conditions by complementing vehicle-end and road-end information [2]. Collaborative vehicle and roadside LiDAR systems hold significant potential, yet their realization faces numerous challenges, especially in the calibration of sensor systems.
Calibration of LiDAR systems across heterogeneous vehicle and infrastructure endpoints is a vital step to ensure the accuracy and consistency of perception system data. This requires high-precision synchronization and registration techniques to integrate LiDAR data from both vehicles and infrastructure [3]. This not only involves complex technical issues such as time synchronization, spatial alignment, and data fusion but also necessitates real-time processing capabilities in dynamic environments.
(a) Vehicle Perspective
(b) Infrastructure Perspective
(c) Vehicle Point Cloud
(d) Infrastructure Point Cloud
(e) Cooperative Perspective
Figure 1: Cooperative Vehicle and Infrastructure LiDAR Calibration Schematic.
To tackle these challenges, the research community has recently seen the emergence of novel approaches. These methods have successfully addressed the issue of time synchronization in collaborative vehicle and infrastructure LiDAR systems [4] [5], ensuring real-time data transmission and the accuracy of data fusion [6]. Despite advancements in spatial alignment, there remain outstanding issues such as the acquisition of spatial initial values [7], the high computational load involved in cross-source point cloud registration [8], or the substantial preliminary costs associated with point cloud mapping techniques [9]. Consequently, achieving spatial alignment in a dynamic environment for vehicle-infrastructure cooperative perception systems while maintaining low computational overhead remains a formidable challenge.
This paper presents a novel calibration method for cooperative vehicle and infrastructure LiDAR systems. The method establishes an affinity matrix for vehicular and roadside nodes based on the Overall IoU metric, enabling the identification of shared matching boxes between vehicle and infrastructure endpoints. Subsequently, it computes and refines the extrinsic parameters associated with matched boxes. The method offers several advantages. Firstly, the proposed method does not require initial extrinsic parameters and meets real-time operation requirements. Secondly, it leverages common target information inherent to traffic scenes, enhancing its generalizability. Furthermore, compared to [10], it exclusively utilizes perception information from target detection, resulting in lower computational complexity and reduced data transmission costs, thereby offering greater potential for practical application. Lastly, V2I-Calib is designed with flexibility, having its components well decoupled, which facilitates easy adaptation to meet specific real-world requirements.
The innovations of this paper include:
  • 1)
    We introduce the Overall IoU metric, which monitors the real-time calibration performance of extrinsic parameters. Central to this, we propose a method for constructing an affinity matrix for vehicle-road nodes, encoding the correlation between vehicular and roadside targets.
  • 2)
    We propose a calibration method for cooperative vehicle and infrastructure LiDAR systems that thoroughly exploits the spatial associations between detection boxes, characterized by its independence from initial extrinsic parameter values and real-time capabilities.
  • 3)
    The effectiveness of our method is validated through comparative and ablation experiments on the DAIR-V2X dataset, which is also classified according to the difficulty levels of the calibration tasks.
The contents of this paper are organized into the following sections. Initially, in Section II, we will provide a detailed review of the existing vehicle-infrastructure LiDAR calibration methods. Subsequently, in Section III, we will elaborate on the calibration method introduced in this paper, along with its constituent modules. Furthermore, in Section IV, we will conduct experiments to validate the effectiveness of our proposed method. Lastly, the paper will conclude with a summary and future outlook in Section V.
II RELATED WORK
Point cloud registration can be categorized into same-source and cross-source source registration . In essence, the calibration challenge for disparate vehicle and infrastructure LiDAR systems broadly constitutes a cross-source point cloud registration issue [11].
While significant advancements have been made in the registration tasks for same-source point clouds [12] [13] [14] , cross-source point cloud registration in vehicle-infrastructure cooperative systems continues to face challenges. Compared to standard point cloud registration, the task is complicated by a combination of factors such as noise, outliers, density variation, partial overlaps, and scale differences. However, compared to other cross-source point cloud registration tasks[8], vehicle-infrastructure cooperative systems benefit from more consistent and stable information inherent in traffic scenarios, such as elements of traffic participation including pedestrians, vehicles, and traffic signs.
Calibration methods for vehicle-infrastructure LiDAR systems can be categorized based on the source of information used during the calibration process: methods based on high-precision maps and those based on perception results.
High-precision map-based calibration methods align point cloud data collected by vehicle-mounted LiDAR with pre-constructed high-accuracy maps to achieve precise spatial calibration [9] [15]. The advantage of these methods is their utilization of rich environmental information in high-precision maps for stable and reliable calibration. However, they are limited by their reliance on high-quality map data, and the updating and maintenance of these maps pose a challenge. Therefore, these methods are most applicable in relatively stable environments, such as urban arteries and highways, where high-precision maps are available.
Perception result-based calibration methods achieve calibration by identifying and matching prominent features in the environment, such as road markings and traffic signs. These are essentially feature-based calibration methods [16][17], adapted with systematic approaches to feature extraction algorithms known as perception algorithms in the context of autonomous driving. These methods often employ multiple perception algorithms [10] to increase robustness through redundancy but may not perform well in terms of real-time processing. Some approaches combine existing perception algorithms with custom-designed feature extraction modules [18], but the general applicability of these methods across diverse scenarios requires further validation. Other methods attempt to incorporate prior information, such as location, to achieve better spatio-temporal calibration [7], yet their effectiveness at larger data scales remains to be tested.
Figure 2: V2I-Calib Workflow. The method starts with two sets of detection boxes obtained from both vehicle and infrastructure ends. We construct an affinity matrix for vehicle-road nodes using Core affinity and Other affinity. Utilizing matching algorithms, we identify common detection boxes that are matched across the vehicle and infrastructure, after which we proceed to calculate and optimize the extrinsic parameters.
III METHODOLOGY
In this study, we introduce a novel vehicle-infrastructure LiDAR calibration method (V2I-Calib), the overall workflow of which is illustrated in Figure 2. The following sections will first define the problem of extrinsic parameter estimation, and then provide a detailed explanation of the proposed method and its individual components.
III-A Problem Formulation
Traditional point cloud registration methods typically seek a rotation matrix $\mathbf{R}$ and translation vector $\mathbf{t}$ that transform one point cloud $\mathbf{P}_{1}$ onto another $\mathbf{P}_{2}$ by minimizing their discrepancy. Mathematically, this is articulated as finding the minimum of an objective function $E(\mathbf{R},\mathbf{t})$, where $E$ is usually a measure of distance between the point clouds:
$\min_{\mathbf{R},\mathbf{t}}E(\mathbf{P}_{1},\mathbf{R}\mathbf{P}_{2}+\mathbf{t})$ (1)
However, in the context of vehicle-infrastructure (V2I) cooperation, the point clouds from the vehicle and infrastructure inherently exhibit significant discrepancies, to which the traditional methods may not be directly applicable. In response, this paper introduces a detection-box-based method. The main concept is to exploit the spatial information of detection boxes shared between the vehicle and infrastructure.
Specifically, let $\mathbf{B}^{(\text{veh})}$ denote the set of detection boxes from the vehicle, and $\mathbf{B}^{(\text{inf})}$ those from the infrastructure. Our objective is to identify the common detection box set $\mathbf{B}^{(\text{c})}$, comprising the common targets detected by both systems. Mathematically, the common detection box set is expressed as:
$\mathcal{B}^{(c)}=\{\mathbf{B}\mid\mathbf{B}\in\mathcal{B}^{(v)}\cap\mathcal{B}^{(i)}\}$ (2)
After extracting these common detection boxes from the original point clouds, we transform them into abstract point clouds $\mathbf{P}^{(\text{c})}$. This transformation process involves mapping the vertices of detection boxes to point sets, which we denote with the function $\text{Vertices}(\cdot)$. Thus, for each common detection box $\mathbf{B}\in\mathbf{B}^{(\text{c})}$, its corresponding abstract point cloud is defined as:
$\mathbf{P}^{(c)}=\bigcup_{\mathbf{B}\in\mathcal{B}^{(c)}}\text{Vertices}(\mathbf{B})$ (3)
Ultimately, the goal is to obtain the optimal extrinsic parameters $\mathbf{T}$ by minimizing the distance between corresponding abstract point clouds:
$\min_{\mathbf{R},\mathbf{t}}E(\mathbf{P}_{1}^{(c)},\mathbf{R}\mathbf{P}_{2}^{(c)}+\mathbf{t})$ (4)
$T=\begin{bmatrix}\mathbf{R}&\mathbf{t}\\ \mathbf{0}^{T}&1\end{bmatrix}$ (5)
By employing this abstract point cloud registration method based on common detection boxes, we can effectively tackle the challenges of cross-source point cloud registration in V2I cooperative scenarios. The crux of the issue now becomes how to identify the common detection box set $\mathbf{B}^{(\text{c})}$ and then fine-tune the extrinsics, which will be elaborated in subsequent chapters.
III-B Affinity Formulation
In this section, we delve into the process of formulating affinity measures, a pivotal step within the vehicle-infrastructure cooperative LiDAR calibration framework. By crafting a meticulously designed affinity function, we are able to effectively evaluate the congruence between vehicle-side and infrastructure-side detection boxes, providing accurate input data for the subsequent matching module. There are two primary challenges in this process: firstly, the vehicle and infrastructure detection boxes are situated in distinct coordinate systems, complicating the direct utilization of their corresponding geometric relationships; secondly, the randomness inherent in cross-source point cloud perceptions makes distinguishing shared and single-end detection objects in real-world scenarios difficult.
Our work ingeniously addresses these issues through the introduction of a hypothesized matching pairs strategy and the oIoU metric. Below, we will describe the formulation of affinity measures in two parts: core affinity and additional affinities.
III-B1 Core Affinity
The idea behind formulating the Core Affinity is to explore the scene matching scores of the available detection box pairs from both the vehicle and infrastructure. To this end, this section introduces the Overall Intersection over Union (oIoU) as a metric to gauge the degree of scene matching. Here is the mathematical expression:
$\footnotesize\emph{oIoU}=\frac{1}{\max(m,n)}\sum_{i=0}^{\mathbf{m}-1}\sum_{j=0}^{\mathbf{n}-1}\emph{IoU}_{3D}(\mathbf{BBox}_{\emph{infra},i},\mathbf{BBox}_{\emph{vehicle},j})$ (6)
$IoU_{3D(}\mathbf{A},\mathbf{B})=\frac{\text{Vol}(\mathbf{A}\cap\mathbf{B})}{\text{Vol}(\mathbf{A}\cup\mathbf{B})}$ (7)
where $m$ and $n$ represent the number of 3d detection boxes detected by the infrastructure and the vehicle, respectively. $\mathbf{BBox}_{\emph{infra},i}$ and $\mathbf{BBox}_{\emph{vehicle},j}$ are the i-th and j-th 3d detection boxes detected by the infrastucture and the vehicle, respectively. $\emph{IoU}_{3D}(\cdot)$ is the 3d intersection over union function, which calculates the volumetric overlap between 3d detection boxes. $\text{Vol}(\cdot)$ denotes the volume of a 3D detection box, and $\cap$ and $\cup$ represent the intersection and union in the 3D space of the detection boxes, respectively. The complexity of the naive traversal algorithm is $O(m*n)$.
The oIoU metric extends beyond the conventional IoU measure, which traditionally concentrates on the correspondence of individual box pairs. The oIoU fully leverages the positional relationships between 3D detection boxes, innovatively providing a comprehensive assessment of the spatial alignment quality within detection scenarios.
Building upon this foundation, the process of constructing Core Affinity, as illustrated in Figure 3, employs the concept of hypothetical matching pairs to define the feasible range of extrinsic parameters search between the vehicle and the infrastructure, as per the method outlined in [19]. The oIoU metric is then applied to assess the quality of spatial alignment for the corresponding extrinsic parameters of these tentative vehicle-infrastructure matching pairs.
Figure 3: Core Affinity Calculation: two 3D detection boxes are selected from the infrastructure end (red) and the vehicle end (green) as assumed matching pairs to compute the extrinsic parameters at that instance. The entire scene is then transformed and merged to calculate the scene affinity using the overlap IoU (oIoU) metric.
III-B2 Additional Affinity Calculation
Beyond the core affinity, this section delves into additional affinity metrics aimed at bolstering the precision and durability of the matching process. One such metric is category affinity, which evaluates the congruence of categories between detection boxes on the vehicle and infrastructure sides in the assumed matching pairs. This consideration is pivotal in diminishing the computational demands associated with core affinity, by ensuring that only detection boxes of matching categories are compared, thereby streamlining the matching process and enhancing its efficiency.
Furthermore, affinities such as size affinity, tracking ID affinity, angle affinity, and length affinity from VIPS[7], as well as appearance affinity from DeepSORT[20], can serve as supplements specific to scenarios to enhance the calibration stability of vehicle-infrastructure LiDAR systems in environments where shared objects are sparse.
In general scenarios, a combination of Core Affinity and Category Affinity proves to be effective, as demonstrated in section IV.
III-C Affinity Matrix Formulation and Matching
The primary objective of this module involves constructing an affinity matrix based on the affinity calculation methods, with the aim of identifying the shared matching pairs among vehicle-infrastructure detection boxes.
Our method views the detection boxes as nodes, building weighted graphs $G_{veh}$ and $G_{inf}$ for the vehicle and infrastructure sides, respectively. At this point, the formulation of the affinity matrix and matching issue can be abstracted as a graph matching problem [21]. We can reformulate it into a quadratic assignment problem [22] , described by the vertex affinity matrix $K_{p}$ and the edge affinity matrix $K_{q}$.However, since the quadratic assignment problem is NP-Hard, its computational complexity can easily become a bottleneck in vehicle-infrastructure collaborative scenarios.
We simplify it by integrating the vertex affinity matrix Kp and the edge affinity matrix Kq, reducing it to a linear assignment problem.For each element $K_{p}(i,j)$ , we consider all edges associated with node $i$ and node $j$. Let $E(i)$ be the set of all edges connected to node $i$, and let $K_{q}(u,v)$ denote the affinity between edges $u$ and $v$. The fusion function $F(\cdot)$ integrates the corresponding elements in $K_{q}$ with $K_{p}(i,j)$, for all $u\in E(i)$ and $v\in E(j)$, resulting in the formation of the affinity matrix $A$. The problem is briefly described by the following formula:
$A(i,j)=\emph{F}\left(K_{p}(i,j),\{K_{q}(u,v)|u\in E(i),v\in E(j)\}\right)$ (8)
After obtaining the affinity matrix that reflects the association degree of vehicle-infrastructure nodes, we can employ the Hungarian matching method [23] to find the optimal match within polynomial time. The transition from a quadratic to a linear matching problem might somewhat diminish the ability to independently adjust various affinities. Nonetheless, this simplification aligns well with the needs of our proposed methodology, which merges core and category affinities. Importantly, it significantly lowers the computational complexity of the matching process.
III-D Extrinsics Resolution and Optimization
The aim of this section is to compute the extrinsic parameters from one or more pairs of matched detection boxes and to ensure the accuracy of these parameters. This process confronts two major challenges: first, the correspondence of the eight point pairs within the matched detection boxes is unknown, as this information is affected by the rotation matrix of the extrinsic parameters; second, the calibration of extrinsic parameters presents itself as an inherently ill-posed issue, necessitating an good initial estimate to strike a balance between accuracy and computational expense.
To address these challenges, let =$\mathcal{B}=\{(\mathbf{B}_{i}^{\text{(inf)}},\mathbf{B}_{i}^{\text{(veh)}})\}_{i=1}^{n}$ denote the set of matched detection box pairs, where $\mathbf{B}_{i}^{\text{(inf)}}$ and $\mathbf{B}_{i}^{\text{(veh)}}$ are the detection boxes from the infrastructure and vehicle, respectively. Based on Equation 3, we derive abstract point clouds $\mathbf{P}^{\text{(inf)}}$ and $\mathbf{P}^{\text{(veh)}}$, which encapsulate the vertex data from matched detection boxes. This transformation redefines the initial cross-source point cloud calibration challenge into a standardized point set registration problem.
Figure 4: Extrinsic Resolution Process. This schematic depicts the stages involved in deriving extrinsic parameters from matched detection boxes. The diagram illustrates the challenges of partial correspondence within the two points lists.
The abstract point clouds exhibit a unique partial correspondence relationship. Specifically, for each $i$, the group created by the bijective mapping between the vertices of $\mathbf{B}_{i}^{\text{(inf)}}$ and $\mathbf{B}_{i}^{\text{(veh)}}$ sets up a group-level correspondence. Nonetheless, the internal correspondence among the point pairs within $\mathcal{C}_{i}$ remains unguaranteed.
$\forall k\in\{8i,\dots,8i+7\},\quad\not\exists\,\mathbf{p}^{\text{(inf)}}_{k}\leftrightarrow\mathbf{p}^{\text{(veh)}}_{k}$ (9)
where $\mathbf{p}^{\text{(inf)}}_{k}$ and $\mathbf{p}^{\text{(veh)}}_{k}$ are the points in $\mathbf{P}^{\text{(inf)}}$ and $\mathbf{P}^{\text{(veh)}}$ respectively, originating from the vertices of the $i$-th matched detection box pair in the dataset.
Based on this characteristic, applying a naive point cloud extrinsic parameter calculation method [19] can provide an rough initial value of the extrinsics. Subsequently, based on this initial value, extrinsic parameter optimization is conducted [24][25] to derive the optimal extrinsic parameter values.
IV EXPERIMENT
In this chapter, we will provide a detailed overview of the experimental design, evaluation metrics, dataset characteristics, and specific experimental settings for assessing the performance of our proposed V2I-Calib method.
IV-A Evaluation Metrics
The core metrics for experimental evaluation include Relative Rotation Error (RRE), Relative Translation Error (RTE), and Success Rate.
Relative Rotation Error (RRE): Measures the accuracy of the rotation component in the calibration result, i.e., the angular difference between the estimated rotation matrix $\mathbf{R}_{e}$ and the ground truth $\mathbf{R}_{t}$
$\theta_{\text{degrees}}=\arccos\left(\frac{\text{tr}(\mathbf{\Delta R})-1}{2}\right)\times\frac{180}{\pi}$ (10)
$\mathbf{\Delta R}=\mathbf{R}_{t}^{-1}\mathbf{R}_{e}$ (11)
Relative Translation Error (RTE): Evaluates the accuracy of the translation vector in the calibration result, i.e., the distance difference between the estimated translation vector $\mathbf{t}_{e}$ and the ground truth translation vector $\mathbf{t}_{t}$.
$\emph{RTE}=||\mathbf{t}_{t}^{-1}-\mathbf{t}_{e}||_{2}$ (12)
Success Rate: Defined as the proportion of successfully completed calibration tasks within a preset error threshold, reflecting the method’s robustness and reliability. Following the criteria in [7], we consider $\emph{RTE}<2m$ as the determinant criterion.
These metrics can provide information on the adaptability and stability of the method across different scenarios.
IV-B Datasets
In this study, we conducted experimental validation using the DAIR-V2X dataset[26], which encompasses a wealth of data collected from Vehicle-Infrastructure Cooperative Autonomous Driving (VICAD) scenarios, including vehicular and infrastructural LiDAR data along with their 3D box annotations. The specifications of the vehicular and roadside LiDAR systems are presented in Table I.
TABLE I: Specifications of LiDAR Equipment
Parameter Roadside LiDAR Vehicle LiDAR
LiDAR Points 300 lines 40 lines
Horizontal Field of View 100° 360°
Max Detection Range 280 meters 200 meters
To enhance the assessment of our method’s effectiveness across varied scenarios, we have instituted a data difficulty categorization. While conventional perception datasets like KITTI have established a difficulty rating based on factors such as the count, dimensions, and levels of occlusion or truncation of objects, similar frameworks for calibration tasks within publicly available datasets are notably absent. Mindful of scenario classifications in extant literature and the absence of established difficulty metrics for calibration, we have methodically classified the DAIR-V2X dataset into distinct levels of difficulty, making the criteria and associated experimental code readily accessible.
Our approach to defining data difficulty draws from the established logic used in perception tasks, factoring in the intricacy of the environment, the spatial relation between vehicles and infrastructure, and the presence of shared objects within the scene. For instance, we utilized categories from the KITTI dataset[27]—such as object count, size, and states of occlusion or truncation—as a basis to evaluate their influence on the calibration process for LiDAR systems between vehicles and infrastructure. It’s important to note that we also accounted for the effect of relative positioning within vehicular-roadside scenarios and the possibility of missing shared objects in counter scenarios, highlighting that a shared zone in point clouds can enable the gathering of required shared objects by broadening the detection range, albeit potentially at the expense of increased detection time.
IV-C Experimental Settings
In this section, a systematic performance evaluation of the proposed method will be conducted through comparative experiments and ablation studies. Contrast experiments are aimed at verifying the superiority of the proposed method by comparing its performance with other methods on datasets of varying difficulty. On the other hand, ablation studies further analyze the effectiveness of individual components/modules. Both experiments were conducted on an experimental platform equipped with an Intel i7-9750H CPU.
IV-C1 Contrast Experiment
The comparative experiments assessed the performance of the method presented in this article against other existing methods on metrics such as Rotation Error (RRE), Translation Error (RTE), and success rate. The existing methods are primarily categorized into traditional approaches represented by VGICP [12] and NDT[25], and perception-based approaches exemplified by VIPS [7] and HPCR-VI [10].
TABLE II: Comparative Results of Different Methods.
Dataset
Difficulty
Methods RRE(°) RTE(m) Success
Rate(%)
Time (s)
easy group VGICP [12] 3.18 6.43 6.0 0.77
NDT [25] 1.88 6.36 6.4 19.82
HPCR-VI [10] 5.15 0.91 81.1 /
VIPS [7] 0.96 0.84 79.3 0.33
V2I-Calib 0.68 0.56 96.8 0.21
hard group VGICP [12] 3.11 6.91 4.1 0.42
NDT [25] 2.33 3.91 6.0 29.62
HPCR-VI [10] 6.43 1.52 43.3 /
VIPS [7] 2.12 2.05 57.5 0.29
V2I-Calib 1.92 1.67 71.8 0.15
As indicated in Table II, it is apparent that traditional point cloud registration algorithms largely fail on the DAIR-V2X dataset, and the existing perception-based methods exhibit significant errors. In contrast, the method proposed in this article achieves better results in terms of both precision and real-time performance. Additionally, by incorporating detection box confidence in the engineering implementation, our method effectively mitigates the problem of calibration precision degradation under multi-target interference compared to [10]. In terms of immediacy, it meets the calibration time requirement of less than 0.35 seconds for common intersection traffic scenarios as analyzed in [18]. Notably, the running times for the methods in the more challenging group are generally lower than those in the simpler group, which is attributable to the fewer shared targets in complex scenes, reducing the computational load.
This series of experimental validations confirms that the proposed vehicle-to-infrastructure calibration method not only provides highly accurate calibration results across various scenarios but also demonstrates its feasibility in terms of immediacy, offering robust support for the practical application of vehicle-to-infrastructure LiDAR systems.
(a) Ideal Scenario
(b) Challenging Scenario
(c) Complex Scenario
(d) Complex Scenario
Figure 5: Comparative Calibration Results Across Diverse Scenarios: (a) Accurate alignment of vehicle and infrastructure detection boxes. (b) Calibration errors due to small-sized detection boxes. (c, d) Good and Suboptimal calibration results combining multiple detection boxes, respectively.
(a) RE Comprison
(b) TE Comprison
(c) Time Cost Comprison
Figure 6: Ablation Comparison of Different Affinity Strategies. The violin plots correspond to the left-side metrics of TE, RE, and Time Cost, while the line charts relate to the success rate metric on the right. We compared four strategies for constructing the affinity matrix based on different combinations of affinities: Strategy 1 combines angle affinity with category affinity, Strategy 2 combines core affinity with category affinity, Strategy 3 employs core affinity alone, and Strategy 4 combines length affinity with category affinity. It is evident that Strategy 2, utilized by our method, surpasses the other strategies in terms of both extrinsic parameter accuracy and computational time.
IV-C2 Ablation experiment
In this section, we aim to validate the effectiveness of the core affinity formulation module presented in Section III-B and the extrinsic parameter optimization module discussed in Section III-D through a series of ablation studies.
TABLE III: Methods Composition Of Different Strategies.
Method Affinity Construction Extrinsic Optimization
V2I-Calib-v3 Length and Angle [7] Yes
V2I-Calib-v2 Core and Category No
V2I-Calib-v1 Core and Category Yes
TABLE IV: Ablation Study On Methods With Different Strategies
Dataset
Difficulty
Methods RRE(°) RTE(m) Success
Rate(%)
Time (s)
easy group V2I-Calib-v3 0.95 0.79 80.3 0.32
V2I-Calib-v2 0.98 0.62 50.0 0.18
V2I-Calib-v1 0.68 0.56 96.8 0.21
hard group V2I-Calib-v3 2.69 2.37 57.2 0.28
V2I-Calib-v2 1.93 1.82 53.4 0.14
V2I-Calib-v1 1.92 1.67 71.8 0.15
Comparing V2I-Calib-v1 with V2I-Calib-v3, we observe that the core affinity formulation module proposed in this paper exhibits enhanced robustness, achieving an approximate 20% increase in the success rate compare. Moreover, we evaluated the impact of the extrinsic parameter optimization module on calibration outcomes. A comparative analysis between V2I-Calib-v1 and V2I-Calib-v2, before and after the introduction of the extrinsic optimization module, reveals that the optimization algorithm significantly improves the precision and success rate of the calibration.
To validate the superiority of the affinity metrics proposed in this paper, we conducted comparative analyses with category affinity, angle affinity, and length affinity as delineated in [7]. The results, as depicted in Figure 6, demonstrate that Strategy 2, which is core affinity, achieves a higher success rate and extrinsic parameter accuracy within an acceptable time frame. Notably, when comparing Strategy 2 (core affinity with category affinity) with Strategy 3 (core affinity without category affinity), it is apparent that the inclusion of category affinity significantly reduces the overall computational time and notably enhances performance within the hard group.
Regarding the Strategy 1 angle affinity and Strategy 4 length affinity extended from [7], these strategies were originally designed for scenarios assisted by positional information. However, experimental results show that they are somewhat applicable in scenarios without initial extrinsic parameter values, although their performance significantly decreases in challenging settings. Upon further analysis of the scenarios where these strategies failed, we observed that most of these scenarios involved extrinsic parameters with relatively large absolute values, aligning with the limitations inherent in their initial design rationale.
These ablation studies not only confirm the efficacy of each distinct module within our framework but also underscore their collective role in enhancing the overall functionality of the system.
V CONCLUSIONS
This study introduces a novel calibration method, V2I-Calib, for cooperative vehicle and infrastructure LiDAR systems. Through a series of experiments, our method demonstrates superior calibration accuracy and robustness across datasets of various difficulties, particularly excelling in handling complex scenes while maintaining real-time performance. In essence, the calibration task for cooperative vehicle-infrastructure LiDAR systems is fundamentally a data association task, bearing some similarity to multi-object tracking tasks.
Future research will integrate the concept of temporal association from multi-object detection, merging spatial alignment and temporal synchronization tasks. Furthermore, the integration of 2D bounding box information will shift the research focus towards multi-sensor fusion calibration in cooperative vehicle-infrastructure systems.
As cooperative vehicle-infrastructure technology continues to evolve, we anticipate this framework to adapt to more application scenarios, providing a solid foundation for the safety and reliability of autonomous driving systems.

### v2

@@
 Abstract
-Cooperative vehicle and infrastructure LiDAR systems hold great potential, yet their implementation faces numerous challenges.
-Calibration of LiDAR systems across heterogeneous vehicle and infrastructure endpoints is a critical step to ensure the accuracy and consistency of perception system data, necessitating calibration methods that are real-time and stable.
-To this end, this paper introduces a novel calibration method for cooperative vehicle and road infrastructure LiDAR systems, which exploits spatial association information between detection boxes.
-The method centers around a novel Overall IoU metric that reflects the correlation of targets between vehicle and infrastructure, enabling real-time monitoring of calibration results.
-We search for common matching boxes between vehicle and infrastructure nodes by constructing an affinity matrix.
-Subsequently, these matching boxes undergo extrinsic parameter computation and optimization.
-Comparative and ablation experiments on the DAIR-V2X dataset confirm the superiority of our method.
-To better reflect the differences in calibration results, we have categorized the calibration tasks on the DAIR-V2X dataset based on their level of difficulty, enriching the dataset’s utility for future research.
-Our project is available at https://github.com/MassimoQu/v2i-calib .
+Cooperative LiDAR systems integrating vehicles and road infrastructure, termed V2I calibration, exhibit substantial potential, yet their deployment encounters numerous challenges.
+A pivotal aspect of ensuring data accuracy and consistency across such systems involves the calibration of LiDAR units across heterogeneous vehicular and infrastructural endpoints.
+This necessitates the development of calibration methods that are both real-time and robust, particularly those that can ensure robust performance in urban canyon scenarios without relying on initial positioning values.
+Accordingly, this paper introduces a novel approach to V2I calibration, leveraging spatial association information among perceived objects.
+Central to this method is the innovative Overall Intersection over Union (oIoU) metric, which quantifies the correlation between targets identified by vehicle and infrastructure systems, thereby facilitating the real-time monitoring of calibration results.
+Our approach involves identifying common targets within the perception results of vehicle and infrastructure LiDAR systems through the construction of an affinity matrix.
+These common targets then form the basis for the calculation and optimization of extrinsic parameters.
+Comparative and ablation studies conducted using the DAIR-V2X dataset substantiate the superiority of our approach.
+For further insights and resources, our project repository is accessible at https://github.com/MassimoQu/v2i-calib.
 I INTRODUCTION
@@
 (e) Cooperative Perspective
-Figure 1: Cooperative Vehicle and Infrastructure LiDAR Calibration Schematic.
+Figure 1: Cooperative V2I LiDAR Calibration Schematic.
 To tackle these challenges, the research community has recently seen the emergence of novel approaches.
-These methods have successfully addressed the issue of time synchronization in collaborative vehicle and infrastructure LiDAR systems [4] [5], ensuring real-time data transmission and the accuracy of data fusion [6].
-Despite advancements in spatial alignment, there remain outstanding issues such as the acquisition of spatial initial values [7], the high computational load involved in cross-source point cloud registration [8], or the substantial preliminary costs associated with point cloud mapping techniques [9].
-Consequently, achieving spatial alignment in a dynamic environment for vehicle-infrastructure cooperative perception systems while maintaining low computational overhead remains a formidable challenge.
-This paper presents a novel calibration method for cooperative vehicle and infrastructure LiDAR systems.
-The method establishes an affinity matrix for vehicular and roadside nodes based on the Overall IoU metric, enabling the identification of shared matching boxes between vehicle and infrastructure endpoints.
+These methods have provided valuable insights into addressing the issue of time synchronization in collaborative vehicle and infrastructure LiDAR systems [4, 5], contributing to the enhancement of real-time data transmission and the accuracy of data fusion [6].
+Despite advancements in spatial alignment, current calibration methods for vehicle-infrastructure cooperative systems typically require initial calibration values as a prerequisite [7, 8], which can be unstable, especially within city canyons.
+Moreover, existing studies [9] have primarily focused on integrating collaborative calibration methods into cooperative perception strategies, which, to some extent, complicates the model’s interpretability.
+This approach makes it challenging to quantify whether spatial alignment issues within the calibration process are adequately addressed, which is crucial for further enhancing the overall perception results.
+A more focused study on LiDAR calibration in cooperative V2X systems should concentrate on the fundamental issues of point cloud registration.
+Registration can be classified into same-source and cross-source types based on the similarity of data collection devices, or into local and global types depending on initial calibration values.
+In essence, the calibration challenge for disparate vehicle and infrastructure LiDAR systems broadly constitutes a cross-source point cloud registration issue [10].
+Furthermore, to maximize the stability of calibration methods within urban canyons, it is advantageous for global point cloud registration to proceed without the necessity for pre-calibration initial values.
+Significant progress has been made in same-source point cloud registration [11, 12], but cross-source registration still faces challenges such as noise, outliers, and variations in density and scale [13].
+These challenges necessitate complex computational solutions.
+Yet, vehicle-infrastructure systems benefit from consistent types of traffic participants in their scenarios, providing a stable foundation for data analysis.
+Advancements in global point cloud registration, like those proposed by [14, 15, 16], offer initial-calibration-free solutions but demand high computational resources, making real-time processing difficult.
+Current research is thus directed towards developing efficient algorithms, like graph-based optimization techniques [17], to reduce computational costs.
+To sum up, existing approaches must balance the provision of initial extrinsic values with computational efficiency.
+This trade-off poses significant challenges for the practical deployment of calibration in V2I cooperative perception systems, necessitating innovative solutions that can bridge the gap between theoretical accuracy and real-world applicability.
+This paper presents a novel calibration method for V2I cooperative LiDAR systems.
+The method establishes an affinity matrix for vehicular and roadside nodes based on the innovative Overall Intersection over Union (oIoU) metric, enabling the identification of shared perception targets between vehicle and infrastructure endpoints.
 Subsequently, it computes and refines the extrinsic parameters associated with matched boxes.
@@
   • 1)
-    We introduce the Overall IoU metric, which monitors the real-time calibration performance of extrinsic parameters.
+    We propose a calibration method for cooperative vehicle and infrastructure LiDAR systems that thoroughly exploits the spatial associations between perception objects, characterized by its independence from initial extrinsic parameter values and real-time capabilities.
+  • 2)
+    We introduce the Overall Intersection over Union metric (oIoU), which monitors the real-time calibration performance of extrinsic parameters.
 Central to this, we propose a method for constructing an affinity matrix for vehicle-road nodes, encoding the correlation between vehicular and roadside targets.
-  • 2)
-    We propose a calibration method for cooperative vehicle and infrastructure LiDAR systems that thoroughly exploits the spatial associations between detection boxes, characterized by its independence from initial extrinsic parameter values and real-time capabilities.
   • 3)
-    The effectiveness of our method is validated through comparative and ablation experiments on the DAIR-V2X dataset, which is also classified according to the difficulty levels of the calibration tasks.
-The contents of this paper are organized into the following sections.
-Initially, in Section II, we will provide a detailed review of the existing vehicle-infrastructure LiDAR calibration methods.
-Subsequently, in Section III, we will elaborate on the calibration method introduced in this paper, along with its constituent modules.
-Furthermore, in Section IV, we will conduct experiments to validate the effectiveness of our proposed method.
-Lastly, the paper will conclude with a summary and future outlook in Section V.
-II RELATED WORK
-Point cloud registration can be categorized into same-source and cross-source source registration .
-In essence, the calibration challenge for disparate vehicle and infrastructure LiDAR systems broadly constitutes a cross-source point cloud registration issue [11].
-While significant advancements have been made in the registration tasks for same-source point clouds [12] [13] [14] , cross-source point cloud registration in vehicle-infrastructure cooperative systems continues to face challenges.
-Compared to standard point cloud registration, the task is complicated by a combination of factors such as noise, outliers, density variation, partial overlaps, and scale differences.
-However, compared to other cross-source point cloud registration tasks[8], vehicle-infrastructure cooperative systems benefit from more consistent and stable information inherent in traffic scenarios, such as elements of traffic participation including pedestrians, vehicles, and traffic signs.
-Calibration methods for vehicle-infrastructure LiDAR systems can be categorized based on the source of information used during the calibration process: methods based on high-precision maps and those based on perception results.
-High-precision map-based calibration methods align point cloud data collected by vehicle-mounted LiDAR with pre-constructed high-accuracy maps to achieve precise spatial calibration [9] [15].
-The advantage of these methods is their utilization of rich environmental information in high-precision maps for stable and reliable calibration.
-However, they are limited by their reliance on high-quality map data, and the updating and maintenance of these maps pose a challenge.
-Therefore, these methods are most applicable in relatively stable environments, such as urban arteries and highways, where high-precision maps are available.
-Perception result-based calibration methods achieve calibration by identifying and matching prominent features in the environment, such as road markings and traffic signs.
-These are essentially feature-based calibration methods [16][17], adapted with systematic approaches to feature extraction algorithms known as perception algorithms in the context of autonomous driving.
-These methods often employ multiple perception algorithms [10] to increase robustness through redundancy but may not perform well in terms of real-time processing.
-Some approaches combine existing perception algorithms with custom-designed feature extraction modules [18], but the general applicability of these methods across diverse scenarios requires further validation.
-Other methods attempt to incorporate prior information, such as location, to achieve better spatio-temporal calibration [7], yet their effectiveness at larger data scales remains to be tested.
+    The effectiveness of our method is validated through comparative and ablation experiments on the DAIR-V2X dataset.
 Figure 2: V2I-Calib Workflow.
-The method starts with two sets of detection boxes obtained from both vehicle and infrastructure ends.
-We construct an affinity matrix for vehicle-road nodes using Core affinity and Other affinity.
-Utilizing matching algorithms, we identify common detection boxes that are matched across the vehicle and infrastructure, after which we proceed to calculate and optimize the extrinsic parameters.
+The process initiates with perception objects from vehicle and infrastructure, forming an affinity matrix using Core and Other affinity.
+Matching algorithms then identify common objects across both ends, facilitating the calculation and optimization of extrinsic parameters.
+II PROBLEM FORMULATION
+Traditional point cloud registration methods typically seek a rotation matrix $\hat{\mathbf{R}}$ and translation vector $\hat{\mathbf{t}}$ that transform one point cloud $\mathbf{P}_{1}$ onto another $\mathbf{P}_{2}$ by minimizing their discrepancy.
+Mathematically, this is articulated as finding the minimum of an objective function $F(\mathbf{P}_{1},\mathbf{P}_{2})$, where $F$ is usually a measure of distance between the point clouds:
+$(\hat{\mathbf{R}},\hat{\mathbf{t}})=\arg\min_{\mathbf{R},\mathbf{t}}F(\mathbf{P}_{1}^{(c)},\mathbf{R}\mathbf{P}_{2}^{(c)}+\mathbf{t})$ (1)
+However, in the context of vehicle-infrastructure (V2I) cooperation, the point clouds from the vehicle and infrastructure inherently exhibit significant discrepancies, to which the traditional methods may not be directly applicable.
+In response, this paper introduces an object-aware method.
+The main concept is to exploit the spatial information of perception objects shared between the vehicle and infrastructure.
+Specifically, let $\mathcal{B}^{(\text{veh})}$ denote the set of perception objects from the vehicle, and $\mathcal{B}^{(\text{inf})}$ those from the infrastructure.
+Our objective is to identify the common perception object set $\mathcal{B}^{(c)}$, comprising the common targets detected by both systems.
+Mathematically, the common detection box set is expressed as:
+$\mathcal{B}^{(c)}=\mathcal{B}^{(veh)}\cap\mathcal{B}^{(inf)}$ (2)
+After extracting common perception objects from the original point cloud, we transform these objects into feature point clouds, denoted as $\mathbf{P}^{(\text{c})}$.
+This transformation involves mapping the vertices of the perception objects into point sets, a process facilitated by the function $\text{Vertices}(\cdot)$.
+Specifically, the vertices of each detection box $\mathcal{B}_{i}$ are extracted and merged to form the complete feature point cloud:
+$\mathbf{P}^{(c)}=\bigcup_{\mathcal{B}_{i}\in\mathcal{B}^{(c)}}\text{Vertices}(\mathcal{B}_{i})$ (3)
+This transformation yields a well-characterized feature point cloud, which can then be processed within the framework of classical point cloud registration algorithms, as formulated in Equation (1).
+Upon obtaining the rotation matrix $\hat{\mathbf{R}}$ and translation vector $\hat{\mathbf{t}}$, these are integrated to produce the optimal extrinsic parameter result $\hat{\mathbf{T}}$.
+$\hat{\mathbf{T}}=\begin{bmatrix}\hat{\mathbf{R}}&\hat{\mathbf{t}}\\ \mathbf{0}^{T}&1\end{bmatrix}$ (4)
+By employing the point cloud registration method based on common perception objects, we can effectively tackle the challenges of cross-source point cloud registration in V2I cooperative scenarios.
+The crux of the issue now becomes how to identify the common detection box set $\mathcal{B}^{(c)}$ and then fine-tune the extrinsics, which will be elaborated in subsequent chapters.
 III METHODOLOGY
 In this study, we introduce a novel vehicle-infrastructure LiDAR calibration method (V2I-Calib), the overall workflow of which is illustrated in Figure 2.
-The following sections will first define the problem of extrinsic parameter estimation, and then provide a detailed explanation of the proposed method and its individual components.
-III-A Problem Formulation
-Traditional point cloud registration methods typically seek a rotation matrix $\mathbf{R}$ and translation vector $\mathbf{t}$ that transform one point cloud $\mathbf{P}_{1}$ onto another $\mathbf{P}_{2}$ by minimizing their discrepancy.
-Mathematically, this is articulated as finding the minimum of an objective function $E(\mathbf{R},\mathbf{t})$, where $E$ is usually a measure of distance between the point clouds:
-$\min_{\mathbf{R},\mathbf{t}}E(\mathbf{P}_{1},\mathbf{R}\mathbf{P}_{2}+\mathbf{t})$ (1)
-However, in the context of vehicle-infrastructure (V2I) cooperation, the point clouds from the vehicle and infrastructure inherently exhibit significant discrepancies, to which the traditional methods may not be directly applicable.
-In response, this paper introduces a detection-box-based method.
-The main concept is to exploit the spatial information of detection boxes shared between the vehicle and infrastructure.
-Specifically, let $\mathbf{B}^{(\text{veh})}$ denote the set of detection boxes from the vehicle, and $\mathbf{B}^{(\text{inf})}$ those from the infrastructure.
-Our objective is to identify the common detection box set $\mathbf{B}^{(\text{c})}$, comprising the common targets detected by both systems.
-Mathematically, the common detection box set is expressed as:
-$\mathcal{B}^{(c)}=\{\mathbf{B}\mid\mathbf{B}\in\mathcal{B}^{(v)}\cap\mathcal{B}^{(i)}\}$ (2)
-After extracting these common detection boxes from the original point clouds, we transform them into abstract point clouds $\mathbf{P}^{(\text{c})}$.
-This transformation process involves mapping the vertices of detection boxes to point sets, which we denote with the function $\text{Vertices}(\cdot)$.
-Thus, for each common detection box $\mathbf{B}\in\mathbf{B}^{(\text{c})}$, its corresponding abstract point cloud is defined as:
-$\mathbf{P}^{(c)}=\bigcup_{\mathbf{B}\in\mathcal{B}^{(c)}}\text{Vertices}(\mathbf{B})$ (3)
-Ultimately, the goal is to obtain the optimal extrinsic parameters $\mathbf{T}$ by minimizing the distance between corresponding abstract point clouds:
-$\min_{\mathbf{R},\mathbf{t}}E(\mathbf{P}_{1}^{(c)},\mathbf{R}\mathbf{P}_{2}^{(c)}+\mathbf{t})$ (4)
-$T=\begin{bmatrix}\mathbf{R}&\mathbf{t}\\ \mathbf{0}^{T}&1\end{bmatrix}$ (5)
-By employing this abstract point cloud registration method based on common detection boxes, we can effectively tackle the challenges of cross-source point cloud registration in V2I cooperative scenarios.
-The crux of the issue now becomes how to identify the common detection box set $\mathbf{B}^{(\text{c})}$ and then fine-tune the extrinsics, which will be elaborated in subsequent chapters.
-III-B Affinity Formulation
-In this section, we delve into the process of formulating affinity measures, a pivotal step within the vehicle-infrastructure cooperative LiDAR calibration framework.
-By crafting a meticulously designed affinity function, we are able to effectively evaluate the congruence between vehicle-side and infrastructure-side detection boxes, providing accurate input data for the subsequent matching module.
-There are two primary challenges in this process: firstly, the vehicle and infrastructure detection boxes are situated in distinct coordinate systems, complicating the direct utilization of their corresponding geometric relationships; secondly, the randomness inherent in cross-source point cloud perceptions makes distinguishing shared and single-end detection objects in real-world scenarios difficult.
-Our work ingeniously addresses these issues through the introduction of a hypothesized matching pairs strategy and the oIoU metric.
-Below, we will describe the formulation of affinity measures in two parts: core affinity and additional affinities.
-III-B1 Core Affinity
-The idea behind formulating the Core Affinity is to explore the scene matching scores of the available detection box pairs from both the vehicle and infrastructure.
-To this end, this section introduces the Overall Intersection over Union (oIoU) as a metric to gauge the degree of scene matching.
-Here is the mathematical expression:
-$\footnotesize\emph{oIoU}=\frac{1}{\max(m,n)}\sum_{i=0}^{\mathbf{m}-1}\sum_{j=0}^{\mathbf{n}-1}\emph{IoU}_{3D}(\mathbf{BBox}_{\emph{infra},i},\mathbf{BBox}_{\emph{vehicle},j})$ (6)
-$IoU_{3D(}\mathbf{A},\mathbf{B})=\frac{\text{Vol}(\mathbf{A}\cap\mathbf{B})}{\text{Vol}(\mathbf{A}\cup\mathbf{B})}$ (7)
-where $m$ and $n$ represent the number of 3d detection boxes detected by the infrastructure and the vehicle, respectively.
-$\mathbf{BBox}_{\emph{infra},i}$ and $\mathbf{BBox}_{\emph{vehicle},j}$ are the i-th and j-th 3d detection boxes detected by the infrastucture and the vehicle, respectively.
-$\emph{IoU}_{3D}(\cdot)$ is the 3d intersection over union function, which calculates the volumetric overlap between 3d detection boxes.
-$\text{Vol}(\cdot)$ denotes the volume of a 3D detection box, and $\cap$ and $\cup$ represent the intersection and union in the 3D space of the detection boxes, respectively.
-The complexity of the naive traversal algorithm is $O(m*n)$.
-The oIoU metric extends beyond the conventional IoU measure, which traditionally concentrates on the correspondence of individual box pairs.
-The oIoU fully leverages the positional relationships between 3D detection boxes, innovatively providing a comprehensive assessment of the spatial alignment quality within detection scenarios.
+III-A Affinity Formulation
+In this section, we explore the development of affinity measures critical to the vehicle-infrastructure cooperative LiDAR calibration framework.
+By creating a detailed affinity function, we effectively assess the alignment between vehicle-side and infrastructure-side perception objects, essential for the next matching module.
+The challenges include differing coordinate systems for perception objects and the randomness in cross-source point cloud data, complicating the identification of shared versus unique detection objects.
+We address these challenges using a hypothesized matching pairs strategy and the Overall Intersection over Union metric(oIoU).
+We’ll discuss the formulation of affinity measures in two parts: core affinity and additional affinities.
+III-A1 Core Affinity
+This section outlines the formulation of Core Affinity by evaluating scene matching scores using the Overall Intersection over Union (oIoU) metric, which measures the extent of scene congruence based on detection box pairs from both vehicle and infrastructure sources:
+$\emph{oIoU}=\frac{1}{\max(m,n)}\sum_{i=0}^{\mathbf{m}-1}\sum_{j=0}^{\mathbf{n}-1}\emph{IoU}_{3D}(\mathcal{B}^{(inf)}_{i},\mathcal{B}^{(veh)}_{j})$ (5)
+$IoU_{3D}(\mathcal{A},\mathcal{B})=\frac{\text{Vol}(\mathcal{A}\cap\mathcal{B})}{\text{Vol}(\mathcal{A}\cup\mathcal{B})}$ (6)
+In this model, $m$ and $n$ are the counts of 3D detection boxes by the infrastructure and vehicle, respectively.
+$\emph{IoU}_{3D}(\cdot)$ calculates volumetric overlaps in 3D space, where $\text{Vol}(\cdot)$ denotes the volumetric overlap between two 3d detection boxes.
+oIoU advances beyond traditional IoU by incorporating positional relationships among boxes, thus providing a comprehensive evaluation of spatial alignment.
 Building upon this foundation, the process of constructing Core Affinity, as illustrated in Figure 3, employs the concept of hypothetical matching pairs to define the feasible range of extrinsic parameters search between the vehicle and the infrastructure, as per the method outlined in [18].
 The oIoU metric is then applied to assess the quality of spatial alignment for the corresponding extrinsic parameters of these tentative vehicle-infrastructure matching pairs.
-Figure 3: Core Affinity Calculation: two 3D detection boxes are selected from the infrastructure end (red) and the vehicle end (green) as assumed matching pairs to compute the extrinsic parameters at that instance.
-The entire scene is then transformed and merged to calculate the scene affinity using the overlap IoU (oIoU) metric.
-III-B2 Additional Affinity Calculation
+Figure 3: SVD-based extrinsic parameters are computed using selected 3D detection boxes from vehicle (green) and infrastructure (red) as assumed matching pairs [18].
+The scene is transformed and merged to assess scene affinity via the oIoU metric.
+III-A2 Additional Affinity Calculation
 Beyond the core affinity, this section delves into additional affinity metrics aimed at bolstering the precision and durability of the matching process.
@@
 In general scenarios, a combination of Core Affinity and Category Affinity proves to be effective, as demonstrated in section IV.
-III-C Affinity Matrix Formulation and Matching
-The primary objective of this module involves constructing an affinity matrix based on the affinity calculation methods, with the aim of identifying the shared matching pairs among vehicle-infrastructure detection boxes.
-Our method views the detection boxes as nodes, building weighted graphs $G_{veh}$ and $G_{inf}$ for the vehicle and infrastructure sides, respectively.
-At this point, the formulation of the affinity matrix and matching issue can be abstracted as a graph matching problem [21].
-We can reformulate it into a quadratic assignment problem [22] , described by the vertex affinity matrix $K_{p}$ and the edge affinity matrix $K_{q}$.However, since the quadratic assignment problem is NP-Hard, its computational complexity can easily become a bottleneck in vehicle-infrastructure collaborative scenarios.
-We simplify it by integrating the vertex affinity matrix Kp and the edge affinity matrix Kq, reducing it to a linear assignment problem.For each element $K_{p}(i,j)$ , we consider all edges associated with node $i$ and node $j$.
-Let $E(i)$ be the set of all edges connected to node $i$, and let $K_{q}(u,v)$ denote the affinity between edges $u$ and $v$.
-The fusion function $F(\cdot)$ integrates the corresponding elements in $K_{q}$ with $K_{p}(i,j)$, for all $u\in E(i)$ and $v\in E(j)$, resulting in the formation of the affinity matrix $A$.
-The problem is briefly described by the following formula:
-$A(i,j)=\emph{F}\left(K_{p}(i,j),\{K_{q}(u,v)|u\in E(i),v\in E(j)\}\right)$ (8)
-After obtaining the affinity matrix that reflects the association degree of vehicle-infrastructure nodes, we can employ the Hungarian matching method [23] to find the optimal match within polynomial time.
-The transition from a quadratic to a linear matching problem might somewhat diminish the ability to independently adjust various affinities.
-Nonetheless, this simplification aligns well with the needs of our proposed methodology, which merges core and category affinities.
-Importantly, it significantly lowers the computational complexity of the matching process.
-III-D Extrinsics Resolution and Optimization
-The aim of this section is to compute the extrinsic parameters from one or more pairs of matched detection boxes and to ensure the accuracy of these parameters.
-This process confronts two major challenges: first, the correspondence of the eight point pairs within the matched detection boxes is unknown, as this information is affected by the rotation matrix of the extrinsic parameters; second, the calibration of extrinsic parameters presents itself as an inherently ill-posed issue, necessitating an good initial estimate to strike a balance between accuracy and computational expense.
-To address these challenges, let =$\mathcal{B}=\{(\mathbf{B}_{i}^{\text{(inf)}},\mathbf{B}_{i}^{\text{(veh)}})\}_{i=1}^{n}$ denote the set of matched detection box pairs, where $\mathbf{B}_{i}^{\text{(inf)}}$ and $\mathbf{B}_{i}^{\text{(veh)}}$ are the detection boxes from the infrastructure and vehicle, respectively.
-Based on Equation 3, we derive abstract point clouds $\mathbf{P}^{\text{(inf)}}$ and $\mathbf{P}^{\text{(veh)}}$, which encapsulate the vertex data from matched detection boxes.
-This transformation redefines the initial cross-source point cloud calibration challenge into a standardized point set registration problem.
-Figure 4: Extrinsic Resolution Process.
-This schematic depicts the stages involved in deriving extrinsic parameters from matched detection boxes.
-The diagram illustrates the challenges of partial correspondence within the two points lists.
-The abstract point clouds exhibit a unique partial correspondence relationship.
-Specifically, for each $i$, the group created by the bijective mapping between the vertices of $\mathbf{B}_{i}^{\text{(inf)}}$ and $\mathbf{B}_{i}^{\text{(veh)}}$ sets up a group-level correspondence.
-Nonetheless, the internal correspondence among the point pairs within $\mathcal{C}_{i}$ remains unguaranteed.
-$\forall k\in\{8i,\dots,8i+7\},\quad\not\exists\,\mathbf{p}^{\text{(inf)}}_{k}\leftrightarrow\mathbf{p}^{\text{(veh)}}_{k}$ (9)
-where $\mathbf{p}^{\text{(inf)}}_{k}$ and $\mathbf{p}^{\text{(veh)}}_{k}$ are the points in $\mathbf{P}^{\text{(inf)}}$ and $\mathbf{P}^{\text{(veh)}}$ respectively, originating from the vertices of the $i$-th matched detection box pair in the dataset.
-Based on this characteristic, applying a naive point cloud extrinsic parameter calculation method [19] can provide an rough initial value of the extrinsics.
-Subsequently, based on this initial value, extrinsic parameter optimization is conducted [24][25] to derive the optimal extrinsic parameter values.
+III-B Affinity Matrix Formulation and Matching
+This module constructs an affinity matrix from vehicle and infrastructure detection boxes, treating them as nodes within weighted graphs $G_{\text{veh}}$ and $G_{\text{inf}}$.
+This setup transforms the matching problem into a graph matching problem [20], often modeled as a quadratic assignment problem (QAP) [21], involving vertex affinity matrix $\mathbf{K}_{p}$ and the edge affinity matrix $\mathbf{K}_{q}$.
+Due to the NP-Hard nature of QAP and its high computational demand, we simplify the problem to a linear assignment format by integrating $\mathbf{K}_{p}$ and $\mathbf{K}_{q}$ into a single affinity matrix $\mathbf{A}$:
+$\mathbf{A}(i,j)=\alpha\cdot\mathbf{K}_{p}(i,j)+\beta\cdot\sum_{u\in E(i),v\in E(j)}\mathbf{K}_{q}(u,v)$ (7)
+Here, $E(i)$ represents the set of all edges connected to node $i$.
+The parameters $\alpha$ and $\beta$ are the adjustment factors for the vertex affinity $\mathbf{K}_{p}(i,j)$ and edge affinity $\mathbf{K}_{q}(i,j)$, respectively, and are empirically set to 0.5.
+The summation aggregates the edge affinities to complement the vertex affinity in the matrix $\mathbf{A}$.
+With affinity matrix $\mathbf{A}$, the Hungarian method [22] efficiently finds optimal matches in polynomial time while reducing computational complexity, suitable for the dynamics of vehicle-infrastructure collaboration.
+III-C Extrinsics Resolution and Optimization
+The objective of this section is to compute extrinsic parameters from matched detection box pairs and to ensure their accuracy, addressing two primary challenges: the unknown correspondence of point pairs within the boxes, influenced by the rotation matrix of the extrinsics, and the ill-posed nature of calibration that requires a good initial estimate to balance accuracy with computational cost.
+After obtaining $\mathcal{B}^{(c)}$, we derive the feature point clouds $\mathbf{P}^{\text{(inf)}}$ and $\mathbf{P}^{\text{(veh)}}$ using Equation (3), transforming the initial cross-source point cloud calibration challenge into a standardized point set registration problem.
+A naive extrinsic parameter calculation method provides an initial value [18], which is optimized to derive the optimal parameters [23, 24].
 IV EXPERIMENT
-In this chapter, we will provide a detailed overview of the experimental design, evaluation metrics, dataset characteristics, and specific experimental settings for assessing the performance of our proposed V2I-Calib method.
 IV-A Evaluation Metrics
@@
 Relative Rotation Error (RRE): Measures the accuracy of the rotation component in the calibration result, i.e., the angular difference between the estimated rotation matrix $\mathbf{R}_{e}$ and the ground truth $\mathbf{R}_{t}$
-$\theta_{\text{degrees}}=\arccos\left(\frac{\text{tr}(\mathbf{\Delta R})-1}{2}\right)\times\frac{180}{\pi}$ (10)
+$\emph{RRE}=\arccos\left(\frac{\text{tr}(\mathbf{\Delta R})-1}{2}\right)$ (8)
 $\mathbf{\Delta R}=\mathbf{R}_{t}^{-1}\mathbf{R}_{e}$ (9)
 Relative Translation Error (RTE): Evaluates the accuracy of the translation vector in the calibration result, i.e., the distance difference between the estimated translation vector $\mathbf{t}_{e}$ and the ground truth translation vector $\mathbf{t}_{t}$.
-$\emph{RTE}=||\mathbf{t}_{t}^{-1}-\mathbf{t}_{e}||_{2}$ (12)
+$\emph{RTE}=||\mathbf{t}_{t}-\mathbf{t}_{e}||_{2}$ (10)
 Success Rate: Defined as the proportion of successfully completed calibration tasks within a preset error threshold, reflecting the method’s robustness and reliability.
@@
 The specifications of the vehicular and roadside LiDAR systems are presented in Table I.
+The outcomes of the experimental scenarios are illustrated in Figure 5.
 TABLE I: Specifications of LiDAR Equipment
@@
 Max Detection Range 280 meters 200 meters
-To enhance the assessment of our method’s effectiveness across varied scenarios, we have instituted a data difficulty categorization.
-While conventional perception datasets like KITTI have established a difficulty rating based on factors such as the count, dimensions, and levels of occlusion or truncation of objects, similar frameworks for calibration tasks within publicly available datasets are notably absent.
-Mindful of scenario classifications in extant literature and the absence of established difficulty metrics for calibration, we have methodically classified the DAIR-V2X dataset into distinct levels of difficulty, making the criteria and associated experimental code readily accessible.
-Our approach to defining data difficulty draws from the established logic used in perception tasks, factoring in the intricacy of the environment, the spatial relation between vehicles and infrastructure, and the presence of shared objects within the scene.
-For instance, we utilized categories from the KITTI dataset[27]—such as object count, size, and states of occlusion or truncation—as a basis to evaluate their influence on the calibration process for LiDAR systems between vehicles and infrastructure.
-It’s important to note that we also accounted for the effect of relative positioning within vehicular-roadside scenarios and the possibility of missing shared objects in counter scenarios, highlighting that a shared zone in point clouds can enable the gathering of required shared objects by broadening the detection range, albeit potentially at the expense of increased detection time.
 IV-C Experimental Settings
-In this section, a systematic performance evaluation of the proposed method will be conducted through comparative experiments and ablation studies.
-Contrast experiments are aimed at verifying the superiority of the proposed method by comparing its performance with other methods on datasets of varying difficulty.
-On the other hand, ablation studies further analyze the effectiveness of individual components/modules.
-Both experiments were conducted on an experimental platform equipped with an Intel i7-9750H CPU.
+This section describes a systematic performance evaluation of the proposed method, which is conducted through comparative experiments and ablation studies.All experiments were performed on an experimental platform equipped with an Intel i7-9750H CPU.
 IV-C1 Contrast Experiment
-The comparative experiments assessed the performance of the method presented in this article against other existing methods on metrics such as Rotation Error (RRE), Translation Error (RTE), and success rate.
-The existing methods are primarily categorized into traditional approaches represented by VGICP [12] and NDT[25], and perception-based approaches exemplified by VIPS [7] and HPCR-VI [10].
+The comparative experiments assessed the performance of the proposed method against other well-performing, global registration methods [15, 16, 17].
 TABLE II: Comparative Results of Different Methods.
@@
 Time (s)
-easy group VGICP [12] 3.18 6.43 6.0 0.77
-NDT [25] 1.88 6.36 6.4 19.82
-HPCR-VI [10] 5.15 0.91 81.1 /
-VIPS [7] 0.96 0.84 79.3 0.33
+easy group FGR [15] 0.98 0.93 59.8 22.23
+Teaser++ [16] 0.99 0.93 60.4 22.74
+Quatro [17] 1.05 1.04 49.5 23.37
 V2I-Calib 0.68 0.56 96.8 0.21
-hard group VGICP [12] 3.11 6.91 4.1 0.42
-NDT [25] 2.33 3.91 6.0 29.62
-HPCR-VI [10] 6.43 1.52 43.3 /
-VIPS [7] 2.12 2.05 57.5 0.29
+hard group FGR [15] 1.15 1.11 24.8 15.17
+Teaser++ [16] 1.17 1.10 25.7 14.79
+Quatro [17] 1.23 1.21 19.2 16.67
 V2I-Calib 1.92 1.67 71.8 0.15
-As indicated in Table II, it is apparent that traditional point cloud registration algorithms largely fail on the DAIR-V2X dataset, and the existing perception-based methods exhibit significant errors.
-In contrast, the method proposed in this article achieves better results in terms of both precision and real-time performance.
-Additionally, by incorporating detection box confidence in the engineering implementation, our method effectively mitigates the problem of calibration precision degradation under multi-target interference compared to [10].
-In terms of immediacy, it meets the calibration time requirement of less than 0.35 seconds for common intersection traffic scenarios as analyzed in [18].
-Notably, the running times for the methods in the more challenging group are generally lower than those in the simpler group, which is attributable to the fewer shared targets in complex scenes, reducing the computational load.
-This series of experimental validations confirms that the proposed vehicle-to-infrastructure calibration method not only provides highly accurate calibration results across various scenarios but also demonstrates its feasibility in terms of immediacy, offering robust support for the practical application of vehicle-to-infrastructure LiDAR systems.
-(a) Ideal Scenario
-(b) Challenging Scenario
-(c) Complex Scenario
-(d) Complex Scenario
-Figure 5: Comparative Calibration Results Across Diverse Scenarios: (a) Accurate alignment of vehicle and infrastructure detection boxes.
-(b) Calibration errors due to small-sized detection boxes.
-(c, d) Good and Suboptimal calibration results combining multiple detection boxes, respectively.
-(a) RE Comprison
-(b) TE Comprison
-(c) Time Cost Comprison
-Figure 6: Ablation Comparison of Different Affinity Strategies.
-The violin plots correspond to the left-side metrics of TE, RE, and Time Cost, while the line charts relate to the success rate metric on the right.
-We compared four strategies for constructing the affinity matrix based on different combinations of affinities: Strategy 1 combines angle affinity with category affinity, Strategy 2 combines core affinity with category affinity, Strategy 3 employs core affinity alone, and Strategy 4 combines length affinity with category affinity.
-It is evident that Strategy 2, utilized by our method, surpasses the other strategies in terms of both extrinsic parameter accuracy and computational time.
+As indicated in Table II, it is evident that the best current global registration methods struggle with the time complexity required for real-time processing when facing large-scale point clouds such as those in the DAIR-V2X dataset.
+Although these methods achieve near 50% success rate and obtain reasonably accurate initial calibration values on the dataset, they lack practical applicability.
+In contrast, the proposed V2I-Calib method significantly outperforms existing global point cloud registration algorithms in terms of runtime and achieves considerably higher accuracy in simple scenarios with lower complexity.
+In hard groups with more occlusions, the precision of our method is impacted due to the lower accuracy of detection results.
+Notably, the running times for the methods in the more challenging group are generally lower than those in the simpler group, which is attributable to the fewer shared targets in hard scenes, reducing the computational load.
 IV-C2 Ablation experiment
-In this section, we aim to validate the effectiveness of the core affinity formulation module presented in Section III-B and the extrinsic parameter optimization module discussed in Section III-D through a series of ablation studies.
+In this section, we aim to validate the effectiveness of the core affinity formulation module presented in Section III-A and the extrinsic parameter optimization module discussed in Section III-C through a series of ablation studies.
 TABLE III: Methods Composition Of Different Strategies.
@@
 V2I-Calib-v1 1.92 1.67 71.8 0.15
-Comparing V2I-Calib-v1 with V2I-Calib-v3, we observe that the core affinity formulation module proposed in this paper exhibits enhanced robustness, achieving an approximate 20% increase in the success rate compare.
-Moreover, we evaluated the impact of the extrinsic parameter optimization module on calibration outcomes.
-A comparative analysis between V2I-Calib-v1 and V2I-Calib-v2, before and after the introduction of the extrinsic optimization module, reveals that the optimization algorithm significantly improves the precision and success rate of the calibration.
-To validate the superiority of the affinity metrics proposed in this paper, we conducted comparative analyses with category affinity, angle affinity, and length affinity as delineated in [7].
-The results, as depicted in Figure 6, demonstrate that Strategy 2, which is core affinity, achieves a higher success rate and extrinsic parameter accuracy within an acceptable time frame.
-Notably, when comparing Strategy 2 (core affinity with category affinity) with Strategy 3 (core affinity without category affinity), it is apparent that the inclusion of category affinity significantly reduces the overall computational time and notably enhances performance within the hard group.
-Regarding the Strategy 1 angle affinity and Strategy 4 length affinity extended from [7], these strategies were originally designed for scenarios assisted by positional information.
+Figure 4: Ablation Comparison of Affinity Strategies.
+Strategy 1 combines core and category affinities; Strategy 2 uses core affinity alone; Strategy 3 combines angle and category affinities; Strategy 4 integrates length with category affinities.
+(a) Ideal Scenario
+(b) Challenging Scenario
+(c) Complex Scenario
+(d) Complex Scenario
+Figure 5: Comparative Calibration Results Across Diverse Scenarios: (a) Accurate alignment of vehicle and infrastructure detection boxes.
+(b) Calibration errors due to small-sized detection boxes.
+(c, d) Good and Suboptimal calibration results combining multiple detection boxes, respectively.
+As indicated in Tables III and IV, the core affinity formulation module introduced in this study significantly enhances robustness, evidenced by a 20% increase in success rates when comparing V2I-Calib-v1 with V2I-Calib-v3.
+Furthermore, the inclusion of the extrinsic parameter optimization module, as evaluated through a comparison of V2I-Calib-v1 and V2I-Calib-v2, substantially improves calibration precision and success rates.
+The superiority of the proposed affinity metrics was validated against category, angle, and length affinities, detailed in [8].
+The results, depicted in Figure 4, show that Strategy 1 (core affinity combined with category affinity) not only achieves higher success rates and accuracy but also reduces computational time significantly compared to Strategy 2 (core affinity alone), especially in challenging scenarios.
+Regarding the Strategy 3 (angle affinity with category affinity) and Strategy 4 (length affinity with category affinity) extended from [8], these strategies were originally designed for scenarios assisted by positional information.
 However, experimental results show that they are somewhat applicable in scenarios without initial extrinsic parameter values, although their performance significantly decreases in challenging settings.
 Upon further analysis of the scenarios where these strategies failed, we observed that most of these scenarios involved extrinsic parameters with relatively large absolute values, aligning with the limitations inherent in their initial design rationale.
-These ablation studies not only confirm the efficacy of each distinct module within our framework but also underscore their collective role in enhancing the overall functionality of the system.
 V CONCLUSIONS
-This study introduces a novel calibration method, V2I-Calib, for cooperative vehicle and infrastructure LiDAR systems.
-Through a series of experiments, our method demonstrates superior calibration accuracy and robustness across datasets of various difficulties, particularly excelling in handling complex scenes while maintaining real-time performance.
-In essence, the calibration task for cooperative vehicle-infrastructure LiDAR systems is fundamentally a data association task, bearing some similarity to multi-object tracking tasks.
-Future research will integrate the concept of temporal association from multi-object detection, merging spatial alignment and temporal synchronization tasks.
-Furthermore, the integration of 2D bounding box information will shift the research focus towards multi-sensor fusion calibration in cooperative vehicle-infrastructure systems.
-As cooperative vehicle-infrastructure technology continues to evolve, we anticipate this framework to adapt to more application scenarios, providing a solid foundation for the safety and reliability of autonomous driving systems.
+This study introduces V2I-Calib, a calibration algorithm tailored for vehicle-infrastructure cooperative LiDAR systems.
+Unlike existing techniques, it does not require initial external parameters, enhancing stability in practical applications.
+Additionally, V2I-Calib satisfies real-time performance criteria, setting it apart from currently effective global point cloud registration methods.
+Future research will focus on integrating temporal synchronization tasks to enhance practicality, alongside incorporating 2D bounding box data to improve multi-sensor fusion calibration.
+As the technology evolves, the V2I-Calib framework is expected to adapt to broader applications, enhancing the safety and reliability of autonomous driving systems.